OpenStudy (anonymous):

Find the Mistake: provided below is a solution to the equation 15g + 10 = 2(7g + 6). Choose the statement that is true about the given solution. Line # 15g + 10 = 2(7g + 6) 1 15g + 10 = 14g + 12 2 g + 10 = 12 3 g = 2 The entire solution is correct. The provided solution is wrong. There is a mistake in Line #1. The provided solution is wrong. There is a mistake in Line #2. The provided solution is wrong. There is a mistake in Line #3

OpenStudy (anonymous):

When g=2 15g + 10 = 2(7g + 6) 15*2+10=2(7*2+6) 40=40 (TRUE)

OpenStudy (anonymous):

So, what do u think?

OpenStudy (anonymous):

Dont u think the solution is correct?